### v3

# Reassessing Noise Augmentation Methods

Abstract
In this study, we investigate if noise-augmented training can concurrently improve adversarial robustness in automatic speech recognition (ASR) systems. We conduct a comparative analysis of the adversarial robustness of four different state-of-the-art ASR architectures, where each of the ASR architectures is trained under three different augmentation conditions: one subject to background noise, speed variations, and reverberations, another subject to speed variations only, and a third without any form of data augmentation. The results demonstrate that noise augmentation not only improves model performance on noisy speech but also the model’s robustness to adversarial attacks.
keywords:
automatic speech recognition, adversarial attacks, adversarial robustness, noise robustness
1 Introduction
Adversarial attacks, also known as evasion attacks, on machine learning models have garnered significant attention due to their potential to compromise the models’ reliability and security [1]. While early research primarily focused on adversarial attacks on computer vision models, there has been a growing interest in the vulnerability of audio processing models, particularly speech processing systems, to such attacks [2, 3, 4, 5]. Concurrently, efforts to enhance noise robustness have been undertaken to improve the usability of speech processing systems in various environments [6]. However, the relationship between noise robustness and adversarial robustness in the audio domain remains largely unexplored. This paper contributes to filling this gap by investigating whether noise-augmented training can also improve adversarial robustness of speech recognition systems.
Noise robustness, also known as environmental robustness [7] reflects a model’s ability to maintain performance despite exposure to various forms of noise and perturbations. Given that real-world audio data often includes inherent noise and environmental disturbances, ensuring noise robustness is essential for reliable audio processing and analysis [8].
Adversarial robustness, on the other hand, refers to a model’s ability to resist adversarial attacks and maintain accurate predictions when faced with manipulated audio samples. These attacks introduce perturbations that are often imperceptible to humans and can deceive the model into making incorrect predictions [2, 3]. In automatic speech recognition (ASR) systems, adversarial attacks can compromise security and privacy, potentially leading to unauthorized access or incorrect processing of sensitive data. For instance, an attacker could alter an audio signal to mislead an ASR system into executing unintended commands, such as purchasing an unwanted product or navigating to a malicious webpage [4]. Thus, adversarial robustness is critical for model security, enabling the detection and prevention of malicious attacks on audio systems [5].
Understanding the interplay between noise robustness and adversarial robustness in the audio domain is crucial [9]. We hypothesize that ASR models with enhanced resistance to acoustic perturbations may inherently exhibit a degree of robustness to adversarial attacks. To test this hypothesis, we conduct a comprehensive empirical analysis, evaluating the impact of noise-augmented training on the adversarial defense capabilities of various state-of-the-art ASR models. Our study encompasses a range of noise types, in particular background disturbances, and reverberations, and assesses the models’ robustness against attacks in the white-box scenario.
2 Background and Related Work
This section provides an overview of adversarial attacks and various types of robustness concerning ASR systems. We also discuss the interdependency between adversarial and noise robustness, drawing on literature from the image domain.
2.1 Adversarial Attacks
Adversarial examples are inputs intentionally modified to induce errors in machine learning models. These perturbations, often imperceptible to humans, can significantly degrade model performance. In speech recognition, adversarial examples exploit ASR system vulnerabilities, resulting in incorrect transcriptions. This study employs a targeted white-box attack to analyze the models’ susceptibility to adversarial attacks.
[2] introduced the first efficient white-box targeted attack against ASR systems, known as the C&W attack. Assuming the attacker has full access to the model, they can create manipulated audio data that transcribes to a specific target phrase. Originally designed for Mozilla’s DeepSpeech model [10], the C&W attack is a gradient-based method that iteratively computes adversarial noise.
The attack minimizes the following loss function with respect to a perturbation $\delta$:
$L(x,\delta,\hat{y})=l_{m}(f(x+\delta),\hat{y})+c\cdot|\delta|_{2}^{2}\enspace,$
where $x$ represents the benign audio signal, $f(\cdot)$ is the ASR model’s function mapping an audio input to the most likely word sequence, $l_{m}$ is the loss function measuring the accuracy of the ASR system’s prediction of the target transcription $\hat{y}$, and $\delta$ is the perturbation constrained to be smaller than a predefined value $\epsilon$. This value is decreased iteratively to minimize distortion. The parameter $c$, which balances the model’s loss against the perturbation, starts small to find an initial adversarial example and then increases to minimize $\delta$.
The resulting adversarial audio files closely resemble the original samples in terms of their dB scale, while leading to a different transcription output by the ASR system. Figure 1 compares the original spectrum of a benign audio clip with its corresponding C&W adversarial counterpart. Note that in line with the definition of adversarial noise, the noise is treated as simply additive.
2.2 Model Robustness
This paper considers two notions of robustness: noise robustness and adversarial robustness. We evaluate how noise robustness, achieved through noise-augmented training, impacts adversarial robustness.
(a) Original audio clip.
(b) C&W adversarial example.
(c) Adversarial noise signal.
Figure 1: The spectrogram of (a) the original audio signal is compared to (b) the spectrogram of its corresponding C&W adversarial example and (c) the spectrogram of the adversarial noise. Transcription of the original audio signal: “but you know more about that than i do sir”. Transcription of the targeted C&W adversarial example: “yes my dear watson i have solved the mystery”.
2.2.1 Noise Robustness
Environmental noise is a common challenge in real-world audio processing, requiring ASR systems to handle disturbances like background chatter, traffic noise, and reverberation. General noise, similar to adversarial noise, can cause non-targeted misclassifications, affecting accuracy and user experience. Strategies to enhance noise robustness include noise augmentation, denoising algorithms, and robust feature extraction techniques [11, 12, 13, 14].
Noise-augmented training is widely recognized as a reliable method for improving noise resilience [7]. This technique integrates noise-perturbed data during neural network training by adding specific noise types to clean data. Leading ASR systems, such as SpeechBrain, use noise augmentation in their training protocols, with the selection and weighting of noise types being crucial for effectiveness [15].
While some approaches focus on architectural modifications to inherently enhance noise robustness, such as changes in hidden layers or input data representation [6], we focus on noise-augmented training without altering model architecture or processing pipelines.
(a) Original audio clip.
(b) Noisy original audio clip.
(c) Noise signal.
Figure 2: The spectrogram of the original audio signal (a) is compared to the spectrogram of its corresponding noisy example, which in this case is constructed by adding background noise to the original audio signal (b) and the spectrogram of the noise (c), which is the difference between the original and its noisy counterpart. Transcription of the original audio signal: ”to their sorrow they were soon undeceived”. Predicted text for the noisy benign example: ”to their sorrow they were assumes under seats”.
2.2.2 Adversarial Robustness
Adversarial robustness concerns a model’s ability to withstand artificially created noise designed to mislead it, as outlined in Section 2.1. Various techniques have been proposed to enhance this robustness, including adversarial training, defensive distillation, gradient masking, regularization techniques, robust optimization, and feature denoising during training [16, 17, 18, 19, 20, 21, 22]. During inference, techniques like input transformations and randomization are employed [23, 24].
Adversarial training involves training the model on both clean and adversarially perturbed samples to enhance its resilience against adversarial manipulations. However, this process is computationally intensive, especially for large datasets. [25] highlighted that while adversarial training is effective for simple datasets, it struggles to scale for high-dimensional data like speech, leaving many blind spots.
This paper evaluates whether noise robustness can enhance adversarial robustness. If successful, this approach could potentially reduce the computational resources needed for generating adversarial examples during training, thereby minimizing blind spots without compromising robustness.
2.2.3 Correlation between Robustness Notions
Noise robustness enables a model to maintain accuracy amid naturally occurring acoustic disturbances, while adversarial robustness refers to resistance against maliciously engineered inputs. The relationship between these robustness types in the image domain is debated: some studies argue for their independence [26, 27], while others suggest a synergistic relationship [28, 29]. The consensus on their interrelation remains elusive.
Conclusions from image-based research do not directly apply to the audio domain [30]. ASR systems possess unique characteristics, and audio adversarial attacks can exploit phenomena like psychoacoustic hiding, absent in visual contexts [4]. While many studies focus on one type of robustness [3, 4, 7], few have examined both in the audio context [31, 32].
Driven by the ongoing debate in the image domain and the lack of comprehensive studies in the audio domain, this research investigates the interrelation between adversarial and noise robustness in ASR models. We evaluate features related to the success of adversarial example generation and the effort required to produce these examples.
3 Methodology
In this section, we describe the methodology used in our experiments.
3.1 ASR Models
We analyze end-to-end ASR systems, which directly transform spoken language into text using a single, unified model, differing from traditional systems with separate stages [33]. We utilize SpeechBrain, an open-source platform for speech technologies [15]. SpeechBrain ASR systems typically consist of three components:
  • Tokenizer: Processes input transcripts to determine corresponding tokens during training and maps output tokens back to text during inference.
  • Acoustic Model: Maps feature vectors derived from the raw input speech into probability distributions over the tokens.
  • Language Model (optional): Enhances prediction accuracy by evaluating the most probable sequences.
We analyze four different architectures from the SpeechBrain toolkit:
For each architecture, we employed three different training regimes:
  • 1 – Baseline (no augmentation): This model serves as a control, trained on a clean dataset without any augmentation to establish standard performance under ideal conditions.
  • 2 – Augmentation with speed variations: Introduces temporal variability in the training data by applying speed perturbations, simulating natural speech tempo variations [41].
  • 3 – Augmentation with speed variations, background noise, and reverberation: Combines speed variations with background noises and reverberations to mimic challenging and realistic acoustic environments [42, 43].
Each model’s architecture and training parameters are detailed in [15].
3.2 Dataset
We use the LibriSpeech dataset, a widely recognized corpus for ASR research that includes thousands of hours of read English speech from audiobooks [44]. It is often used as a benchmark for evaluating speech recognition models. Our models are trained on the first 100 hours of the clean LibriSpeech dataset. Note that some pre-trained models, like wav2vec, may have been exposed to larger portions of LibriSpeech.
For speech augmentation, we apply several techniques to artificially corrupt the clean audio signals provided by SpeechBrain. These techniques include:
  • Reverberation: Applied to 10% of the files to simulate echo effects.
  • Background Noise: Random samples from the Freesound portion of the MUSAN corpus, which include 843 recordings of music, speech, and background noises [7]. The recordings are incorporated into the clean signal at random signal-to-noise ratios.
  • Speed Variations: The audio signal is re-sampled at three different speeds: reduced by 5%, original rate, and increased by 5%.
These augmentation techniques help simulate real-world acoustic environments and assess the robustness of ASR systems under varying conditions. Figure 2 illustrates the spectrum of a benign audio clip and its corresponding noisy counterpart.
3.3 Adversarial Attacks
In our experiments, we utilize the C&W attack, a prominent adversarial technique initially designed to target the DeepSpeech ASR system. The implementation by [31] for SpeechBrain enables us to evaluate the robustness of different SpeechBrain-based models. We generate 300 adversarial examples per architecture per model using the C&W attack.
For the C&W attack, we base our implementation on the robust_speech framework [31], adhering to the parameters defined therein for consistency and comparability, with the maximum number of iterations set to 3000. This allow us to systematically assess the vulnerability of SpeechBrain models to sophisticated adversarial threats.
3.4 Statistical Tests
To assess whether adversarial samples generated for two different training regimes have the same underlying distribution, we apply the Kolmogorov-Smirnov (KS) test. We use a significance level of $\alpha=0.05$. For $p\leq 0.05$, any KS statistic above $0.1$ indicates a statistically significant difference.
4 Experiments
This section outlines the experiments conducted to evaluate the robustness of ASR models against noise and adversarial attacks.
4.1 Evaluating Noise Robustness
In our initial experiment, we assess the effectiveness of noise-augmented training on enhancing noise robustness in SpeechBrain models. For each architecture, we verify whether this training method improves the models’ performance against natural acoustic perturbations, as described in subsection 3.2. The evaluation involves two metrics:
  • Word Error Rate (WER): WER is a standard ASR performance metric that compares the recognized word sequence with the reference text. It is defined as:
    $\text{WER}=100\cdot\frac{S+D+I}{N}\enspace,$ (1)
    where $S$, $D$, and $I$ represent the numbers of substituted, deleted, and inserted words, respectively, and $N$ denotes the total word count of the reference text.
  • Success Rate: The success rate is the percentage of samples correctly recognized as the given reference text. It is calculated as:
[…]
    where $N_{S}$ is the number of audio clips with no transcription error compared to the reference text, out of a total of $N_{T}$ examples. For benign samples, the ground truth serves as the reference, whereas for C&W adversarial examples, we use the target transcription.
4.2 Evaluating Adversarial Robustness
We employ two distinct types of distortion measurements to quantify the imperceptibility of adversarial perturbations, offering insights into the audibility of the attacks. Let $x(t)$ denote the original continuous audio file and $X(k)$ denote the discretized version:
[…]
    $\text{dB}(x)=20\cdot\max_{t}\log_{10}x(t)\text{.}$
    Subsequently, the distortion level is determined as the difference between the relative loudness of the perturbation $\delta$ and the audio signal $x$, defined as:
    $\text{dB}_{\text{x}}(\delta)=\text{dB}(x)-\text{dB}(\delta)\text{,}$
    assuming equal lengths of $x$ and $\delta$.
[…]
    $\text{SNR${}_{\text{Seg}}$}=\frac{10}{M}\cdot\sum_{m=0}^{M-1}\log_{10}\frac{\sum_{k=mF}^{mF+F-1}X(k)^{2}}{\sum_{k=mF}^{mF+F-1}\delta(k)^{2}}\text{,}\enspace$
    where $M$ represents the number of frames in a signal, $F$ is the frame length, $X$ denotes the clean audio signal, and $\delta$ the adversarial perturbation, assuming equal lengths of $X$ and $\delta$.
[…]
5 Results
In the following we present and discuss the results of our experiments. A demo is available at https://matiuste.github.io/SPSC_24/.
Benign Data Noisy Data
Model Augm. Succ. rate $\uparrow$ WER $\downarrow$ Succ. rate $\uparrow$ WER $\downarrow$ $\Delta$ WER
Wav2vec 1 85.67% 3.11% 81.33% 5.39% 2.28%
2 87.00% 2.84% 83.00% 5.06% 2.22%
3 86.33% 2.95% 82.33% 5.12% 2.17%
[…]
2 71.33% 8.06% 55.67% 20.02% 11.96%
3 71.00% 8.57% 65.33% 11.51% 2.95%
[…]
2 71.33% 8.18% 54.67% 20.02% 11.85%
3 66.67% 10.01% 56.67% 14.35% 4.34%
[…]
2 66.67% 9.29% 53.00% 19.47% 10.18%
3 68.00% 8.51% 53.67% 18.85% 10.34%
Table 1: Performance of the considered ASR systems on benign and noisy data, averaged over 300 utterances. WER and success rate are measured w.r.t. the ground truth transcription. The arrows indicate the direction for models that are more noise robust, e.g., models with a higher success rate are considered more robust against noisy examples. The bold numbers indicate the most robust system per model with respect to the chosen parameter.
C&W Adversarial Data
[…]
Wav2vec 1 100.0% 0.00 155 35.41 18.54
2 99.6% 0.05 343 29.59 14.08
3 100.0% 0.00 158 35.40 18.28
[…]
2 86.0% 7.16 323 27.46 9.47
3 74.0% 14.89 479 23.92 6.69
[…]
2 99.6% 0.05 113 41.13 18.84
3 96.0% 1.53 186 37.65 15.93
[…]
2 100.0% 0.00 18 54.52 27.23
3 100.0% 0.00 20 54.90 27.59
Table 2: Results for C&W targeted attack. All values are averaged over 300 adversarial examples. WER and success rate are measured w.r.t. the target adversarial transcription. The arrows indicate the direction for models that are more robust against adversarial examples, e.g., models with a lower success rate are considered more robust against adversarial examples. The bold numbers indicate the more robust model with respect to the chosen parameter. Underlined numbers show a statistical significant difference between training regime 1 and the considered training regime according to the KS test.
5.1 Evaluating Noise Robustness
In evaluating noise robustness, the reference text for WER and success rate calculations is the ground-truth transcript of the clean test sample. We aim for ASR models to exhibit low WER and high success rate, indicating accurate recognition of the ground-truth transcription. This step is essential to validate the effectiveness of our noise-augmented training regimen before examining the relationship between noise and adversarial robustness.
Our experimental results corroborate that models subjected to noise-augmented training exhibit enhanced robustness when confronted with speech samples with environmental noise. For each model considered in Table 1, the success rate and WER for noisy data have improved with noise-augmented training (the models marked in blue or red). This is evidenced by a lower WER in comparison to their counterparts trained without any augmentation (marked in black).
This improvement is particularly pronounced for the seq2seq models, which build on minimal pre-trained components, while the enhancement is less apparent for the wav2vec and transformer models that rely more on pre-trained parts, especially for feature extraction. Interestingly, the wav2vec-based models maintain a high success rate even with noisy data, whereas for all other architectures, the success rate drops significantly. This could be attributed to the fact that the wav2vec models incorporate pre-trained models that were exposed to noisy data during training. These findings indicate that noise-augmented training indeed enhances noise robustness, although the extent of improvement varies depending on the architecture.
In the next experiment, we evaluate if it also enhances adversarial robustness.
5.2 Evaluating Adversarial Robustness
In the subsequent experiment, we focus on assessing adversarial robustness. For every C&W adversarial example we gather data similar to [46]. To gauge the severity of performance degradation, we compare the success rate of adversarial attacks against our models (see Eq. 2) and analyze the WER (see Eq. 1). Additionally, we evaluate the computational complexity of generating adversarial examples by recording the first hit, defined as the number of iterations required to craft the first successful adversarial example.
Table 2 encapsulates the success rate, computational effort, and perceptibility metrics across various architectures and models for the targeted C&W attack. Our findings reveal that for the seq2seq models trained from scratch, those that underwent noise-augmented training demonstrate a slightly reduced adversarial success rate, a higher WER, or more perceptible noise, indicating improved robustness to adversarial attacks. Again, the effects on improved adversarial robustness are most apparent for the seq2seq-based models. Similarly to the noise robustness, only minor improvements can be seen for the wav2vec- and transformer-based models. Interestingly, we did not find a clear connection between the model’s WER on the benign data (see Table 1, fourth column) and the adversarial success rate (see Table 2, third column).
Our findings indicate that incorporating speed or noise augmentation into the training process enhances the adversarial robustness of most models when confronted with targeted white-box attacks. In cases where data augmentation does not lead to higher robustness compared to the base model, it also does not result in a severe robustness reduction. This underlines that noise augmentation should be used as a standard routine in the training of neural network-based ASR systems.
6 Discussion and Conclusion
Our comprehensive study has shed light on the efficacy of noise augmentation as a strategy for improving the adversarial robustness of automatic speech recognition (ASR) systems. By conducting a systematic comparative analysis of four state-of-the-art ASR architectures, each trained with different augmentation techniques and subjected to targeted white-box attacks, we have provided empirical evidence supporting the benefits of noise augmentation.
Our findings reveal that across the board, ASR models trained with noise augmentation – encompassing background noise, speed variations – and reverberations exhibit enhanced robustness against adversarial attacks compared to those without such augmentation. Interestingly, the implementation of speed variations alone as a form of augmentation also contributed positively to the robustness of the models, albeit to a lesser extent than the combination of multiple noise types. This hints that robustness gains may be attributed to the increased diversity and volume of the dataset. Through our investigation, it becomes evident that incorporating speed and noise augmentation into the training pipeline is beneficial for both noise robustness and adversarial robustness.
The findings from our study suggest several intriguing directions for further research. First, while noise augmentation has proven effective, understanding the optimal combination and intensity of different augmentation types remains an open question. The relative contributions of background noise, speed variations, and reverberations need to be quantified to develop more targeted augmentation strategies. Additionally, the role of dataset size and diversity in enhancing robustness merits deeper exploration. Our results indicate that simply increasing the dataset size with varied augmentations can lead to significant robustness improvements. Future studies could systematically vary the size and diversity of training data to better understand this relationship. Moreover, our research highlights the complex interplay between different forms of augmentation and their impact on adversarial robustness. A thorough investigation into how these augmentations interact with various attack types could provide more granular insights into defense mechanisms.
For future work, we believe that exploring the interplay between noise augmentation and other defense strategies, such as adversarial training and model ensembling, could yield a more robust and multi-faceted approach to securing ASR systems against a wide array of threats. Combining these strategies could leverage the strengths of each, potentially leading to ASR systems that are not only robust against noise and adversarial attacks but also adaptable to evolving threats in real-world applications. By continuing to refine augmentation techniques and exploring their integration with other defensive measures, we can move towards developing ASR systems that are both resilient and reliable in the face of diverse and sophisticated adversarial challenges.
7 Acknowledgements
This work was partly supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2092 CASA – 390781972 and the Wilhelm and Günter Esser Foundation.